OpenStudy (anonymous):

The external magnetic field in a loop decreases from 0.1 Tesla to zero in 2 seconds. The loop of wire has ten turns, a radius of 10 cm, with its axis 45 degrees relative to the magnetic field. How large is the induced voltage in the loop? A. 0 mV B. 11 mV C. 9.4 mV D. 32.5 mV E. 0.31 mV

OpenStudy (anonymous):

The induced voltage is given as "rate of change of magnetic flux". Flux is given as \[\Phi = NAB _{p}\] N = no of turns in the coil A= area of the coil "Bp" = magnitude of magnetic field perpendicular to the area Initially Bp = 0.1 cos 45 Finally Bp = 0 Therefore, initial flux = N*A* 0.1*cos(45) final flux = 0 So change in flux = N*A*0.1*cos(45) - 0 = N*A*0.1*cos(45) Time taken for this change to happen = 2seconds. Voltage =change in flux/time taken =N*A*0.1*cos(45) / 2 Put the value of N and A and find the voltage.